# A Banach *-algebra associated with the action of a group on a topological space

### v1

@@ Abstract.
Given a discrete and countable infinite group $G$ acting via homeomorphism on a compact and Hausdorff space $X$, we consider $\Sigma$ the dynamical system associated to the action. One can naturally associate to $\Sigma$ the crossed product type Banach *-algebra $\ell^{1}(\Sigma)$. We will study how different dynamical properties of $\Sigma$ are characterized as analytic-algebraic properties of $\ell^{1}(\Sigma)$. The dynamical properties of $\Sigma$ that we will study are topological freeness, minimality, topological transitivity and residual topological freeness. We will also see how, under some assumptions on $G$, one can detect if $\Sigma$ is free by detecting closed non-self-adjoint ideals of $\ell^{1}(\Sigma)$.
@@ 1. Introduction
The crossed product is a usual construction in the theory of operator algebras that allows one to obtain an algebra arising for the action of a group on a $C^{*}$-algebra. It is desirable that the crossed product has a Banach algebra structure, and the most common way to do that is by endowing the algebraic skeleton of the algebra with a $C^{*}$-seminorm in order to take the completion of the algebra, which will be a $C^{*}$-algebra. These $C^{*}$-crossed products have been studied at length through the years, and are a really important family of examples of $C^{*}$-algebras.
@@
A more concrete family of examples comes from the action of a countable infinite discrete group $G$ on a compact Hausdorff space $X$ via homeomorphisms. We denote $\Sigma$ the dynamical system associated with this action. In this situation, the action of the group on the space gives rise to an action on the algebra of continuous functions $C(X)$, which is a commutative $C^{*}$-algebra, and from this context one can construct a crossed product endowed with a $C^{*}$-norm. Nevertheless, there are many other types of Banach algebras that can be obtained from the action of a group on a $C^{*}$-algebra, and in the latter years there has been a rise of interest in those. For example, the theory of $L^{p}$-operator algebras has been sparkling in recent years, and an analogue $L^{p}$-operator algebra crossed product can be constructed; see, for example the survey paper [Gar21] by Gardella. One can also consider the involutive Banach algebra $\ell^{1}(\Sigma)$, a concrete example whose $C^{*}$-envelope is the usual $C^{*}$-crossed product. In this paper we are interested in the study of this algebra. This algebra has a very rich structure; for example, it can have closed non-self-adjoint ideals, which is something that cannot occur in its $C^{*}$-envelope because, in a $C^{*}$-algebra, every closed ideal is self-adjoint. This will be of interest to us, as we will see that the existence of closed non-self-adjoint ideals is, to some extent, related to the action not being free.
For actions of $\mathbb{Z}$, $\ell^{1}(\Sigma)$ has been studied by de Jeu, Svensson and Tomiyama in [dJST12], and by de Jeu and Tomiyama in [dJT12, dJT13, dJT16, dJT18]. In [dJST12], it was shown how different dynamical properties are related to analytic-algebraic properties of the ideal structure of $\ell^{1}(\Sigma)$; the results proved there are the main inspiration for this paper. We will generalize some of these theorems to the context of countable infinite discrete groups, as we will shortly describe. We will sometimes have to add the hypothesis of topological freeness of $\Sigma$ (see Definition 3.2), that in many results of [dJST12] was not assumed because, in the context of $\mathbb{Z}$-actions, it may be automatic given some other assumptions. The need to assume topological freeness in the context of arbitrary countable discrete groups does not come as a surprise, given that it was mandatory in the context of $C^{*}$-algebraic crossed products; see, for instance, [AS94, Theorem 2].
We will state the main results of this paper. In what follows, we will assume that $G$ is a countable infinite discrete group and that $X$ is compact Hausdorff space. All ideals are assumed to be two-sided. We will not assume that ideals are neither closed nor self-adjoint.
@@
    Every closed nonzero ideal of $\ell^{1}(\Sigma)$ has nonzero intersection with $C(X)$.
@@
    Every closed and self-adjoint nonzero ideal of $\ell^{1}(\Sigma)$ has nonzero intersection with $C(X)$.
@@
    $\Sigma$ is topologically free.
@@
Equivalence between (1) and (3) can actually be seen as an interpretation for $\ell^{1}(\Sigma)$ of [BK24, Theorem 5.9], which is a result on the more general context of twisted Banach algebra crossed products. However, the proof here is in the spirit of [dJST12, Theorem 4.1], of which it is a generalization for arbitrary discrete and countable (infinite) groups. We point the distinction between closed and self-adjoint ideals and just closed ideals done in (1) and (2): as we mentioned before, in this context it is not automatic that every closed ideal is self-adjoint, so the difference is meaningful. This distinction will be recurrent later in the paper.
@@
Theorem A will prove to be an extremely useful tool. Assuming topological freeness of $\Sigma$, it will allow us to easily see how different dynamical properties of $\Sigma$ are characterized as analytic-algebraic properties of $\ell^{1}(\Sigma)$. The dynamical properties we will characterize are minimality, topological transitivity and residual topological freeness (Definitions 3.11, 3.14 and 3.17, respectively); see Theorems 3.13, 3.16 and 3.22. The characterizations of minimality and topological transitivity generalize [dJST12, Theorems 4.2, 4.10], which were proved for $\mathbb{Z}$-actions. In these results, topological freeness of the action is not assumed, because for $\mathbb{Z}$-actions that is automatic given the minimality or topological transitivity.
@@
An interesting question is how much information of the action $\ell^{1}(\Sigma)$ retains. There are precedents that show that algebras that are of type $L^{1}$ retain more information that their $C^{*}$ counterparts. A great example of that comes as Wendel’s theorem (see [Wen51]). This theorem says that two locally compact Hausdorff groups are isomorphic (as topological groups) if and only if their $L^{1}$-group algebras are isometrically isomorphic. This means that $L^{1}$-group algebras retain all the information of the group, something that the $C^{*}$-group algebras do not satisfy as, for example, the group $C^{*}$-algebras $C^{*}(\mathbb{Z}_{4})$ and $C^{*}(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$ are isomorphic but $\mathbb{Z}_{4}$ is not isomorphic to $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$  [^1] In this paper, given $n\in\mathbb{N}$, we refer to $\mathbb{Z}/n\mathbb{Z}$ (the integers modulo $n$) as $\mathbb{Z}_{n}$..
It is clear that all the information of $\Sigma$ retained by its associated $C^{*}$-crossed product is also retained by $\ell^{1}(\Sigma)$, as the $C^{*}$-crossed product is the envelope of $\ell^{1}(\Sigma)$, but it would be interesting to identify extra dynamical information that is carried by $\ell^{1}(\Sigma)$. In this vein, in [dJST12, Theorem 4.4] it was proved for $\mathbb{Z}$-actions that $\Sigma$ is free (i.e., the action of every element of G apart from the unit has no fixed points) if and only if every closed ideal of $\ell^{1}(\Sigma)$ is self-adjoint. This means that, for $\mathbb{Z}$-actions, $\ell^{1}(\Sigma)$ detects freeness of the action, something that the $C^{*}$-algebraic crossed product does not, as in any $C^{*}$-algebra every closed ideal is automatically self-adjoint.
We cannot generalize [dJST12, Theorem 4.4] as it is because of various reasons. First of all, the proof uses the existence of a closed non-self-adjoint ideal in $\ell^{1}(\mathbb{Z})$, but that is not necessarily true for an arbitrary countable discrete group. It also makes use of the fact that, for $\mathbb{Z}$-actions, the action not being free is equivalent to having a finite orbit, something that is not true in general for other groups. At last, it also subtly uses the fact that every nontrivial subgroup of $\mathbb{Z}$ is isomorphic to $\mathbb{Z}$ (in particular, any nontrivial point-stabilizer subgroup would be isomorphic to $\mathbb{Z}$).
Actually, the statement in [dJST12, Theorem 4.4] does not generalize to nonabelian groups (see Example 4.1). Regardless, it can be shown that freeness of the action always implies that every closed ideal is self-adjoint (see Corollary 4.2). In order to prove the other implication, we have to impose some conditions on the group, but we can manage the following concrete statement for abelian groups.
Theorem B.
Suppose that $G$ is an abelian torsion-free group. Then, $\Sigma$ is free if and only if every closed ideal of $\ell^{1}(\Sigma)$ is self-adjoint.
@@
We will briefly describe the structure of the paper. In Section 2 we will explain the context in which we will be working, we will introduce the basic definitions we need and fix the notation we will use. In particular, we will introduce $\ell^{1}(\Sigma)$, our main object of study. In Section 3, we will do some technical work in order to prove Theorem A. The characterizations of minimality, topological transitivity and residual topological freeness will follow from Theorem A. In Section 4, we will study how to relate non-freeness of the action with the existence of closed non-self-adjoint ideals in $\ell^{1}(\Sigma)$. We will first prove that if the action is free, then every closed ideal of $\ell^{1}(\Sigma)$ is self-adjoint. Finally, we will show how the existence of a nontrivial point-stabilizer subgroup satisfying some conditions implies the existence of a closed non-self-adjoint ideal in $\ell^{1}(\Sigma)$ (see Proposition 4.5); this result implies Theorem B.
@@
Acknowledgments: The author is deeply thankful to his late advisor Fernando Abadie, whose guidance was essential to make all this work possible.
@@ 2. Definitions and preliminaries
In this section we will introduce the basic objects playing a role in this work, and we will fix the notation that we will use. In what follows, $G$ is a countable infinite discrete group and $X$ is a compact Hausdorff space. Given a subset $Y$ of $X$, we denote $\overline{Y}$ its closure in $X$. We denote $e$ the identity element of $G$.
2.1. The algebra $\ell^{1}(\Sigma)$
Let $G$ act by homeomorphisms on $X$. For every $s\in G$, we denote $\sigma_{s}:X\to X$ the homeomorphism associated with the action of $s$ in $X$. We call $\Sigma=(G,X,\sigma)$ a dynamical system. We denote $e$ the unit of $G$, and we denote $C(X)=\{f:X\to\mathbb{C};f\text{ is continuous}\}$ the $C^{*}$-algebra of continuous functions on X. The action of $G$ on $X$ induces an action of $G$ on $C(X)$, that we denote $\alpha:G\to\text{Aut}(C(X))$, given by
@@ 2.1. The algebra $\ell^{1}(\Sigma)$
for every $s\in G,f\in C(X)$ and $x\in X$. It is clear that for every $s\in G$ the map $\alpha_{s}:C(X)\to C(X)$ is an automorphism that preserves the involution in $C(X)$, i.e, $\alpha_{s}(\overline{f})=\overline{\alpha_{s}(f)}$ for every $f\in C(X)$.
@@
$\ell^{1}(\Sigma)=\left\{f:G\to C(X):\sum_{s\in G}\lVert f(s)\rVert_{\infty}<\infty\right\},$
where $\lVert\>\cdot\>\rVert_{\infty}$ denotes the supreme norm in $C(X)$. If we set $\lVert f\rVert=\sum_{s\in G}\lVert f(s)\rVert_{\infty}$ for every $f\in\ell^{1}(\Sigma)$ then $\ell^{1}(\Sigma)$ with the pointwise sum and product is a Banach space. We endow $\ell^{1}(\Sigma)$ with the following product: given $f,g\in\ell^{1}(\Sigma)$ we define
$(fg)(s)=\sum_{j\in G}f(j)\alpha_{j}(g(j^{-1}s))\text{ for every }s\in G.$
We also define an involution $*:\ell^{1}(\Sigma)\to\ell^{1}(\Sigma)$ the following way: given $f\in\ell^{1}(\Sigma)$, we define
@@
With these product and involution we have that $\ell^{1}(\Sigma)$ is a Banach *-algebra. We can canonically embed $C(X)$ in $\ell^{1}(\Sigma)$: given $g\in C(X)$, we identify it with the element of $\ell^{1}(\Sigma)$ that maps $e$ to $g$ and every other element of $G$ to the zero function. This mapping from $C(X)$ to $\ell^{1}(\Sigma)$ is an isometry. Given $g\in C(X)$ we also denote by $g$ to the corresponding element in $\ell^{1}(\Sigma)$.
We will now describe a useful way to work with the elements of $\ell^{1}(\Sigma)$. For every $s\in G$, we define $\delta^{s}\in\ell^{1}(\Sigma)$ as:
$\delta^{s}(s^{\prime})=\begin{dcases}1&\text{if }s^{\prime}=s,\\ 0&\text{if }s^{\prime}\neq s,\\ \end{dcases}$
where $1$ denotes the identity element of $C(X)$ with respect to multiplication, i.e., the function taking constant value $1$. With this notation, we can describe an element $f\in\ell^{1}(\Sigma)$ as:
$f=\sum_{s\in G}f(s)\delta^{s}=\sum_{s\in G}f_{s}\delta^{s},$
where $f_{s}$ denotes $f(s)\in C(X)$ for every $s\in G$. We will usually use this description of the elements of $\ell^{1}(\Sigma)$ without explicit mention. With this notation, we have that
$(f_{i}\delta^{i})(g_{j}\delta^{j})=f_{i}\alpha_{i}(g_{j})\delta^{ij}$
for every $i,j\in G$ and $f_{i},g_{j}\in C(X)$. Also, with this notation we have that the involution is given by the rule
$(f_{s}\delta^{s})^{*}=\alpha_{s}^{-1}(\overline{f_{s}})\delta^{s^{-1}}\text{ for every }s\in G.$
We note that, for every $s\in G$, the elements $\delta^{s}$ are unitary elements of $\ell^{1}(\Sigma)$.
@@
We define the canonical projection $E:\ell^{1}(\Sigma)\to C(X)$ as the map given by $E(f)=f(e)$ for every $f\in C(X)$.
We mention that one can take the $C^{*}$-envelope of $\ell^{1}(\Sigma)$ in order to obtain a $C^{*}$-algebraic crossed product. There are two ways of doing that. Through one of the ways we get the universal crossed product $C^{*}$-algebra associated to $\Sigma$, that is usually denoted $C^{*}(\Sigma)$. By the other we get the reduced crossed product $C^{*}$-algebra associated to $\Sigma$, which is usually denoted as $C^{*}_{r}(\Sigma)$. In both cases we have that $\ell^{1}(\Sigma)$ is a dense *-subalgebra of the $C^{*}$-crossed product. We will not work with these algebras in this paper, but we might mention them in order to compare the situation between $\ell^{1}(\Sigma)$ and their $C^{*}$-counterparts.
We will briefly give some definitions related to the dynamical system $\Sigma$. Given $x\in X$, we denote $G_{x}$ the stabilizer subgroup of $x$ as the subgroup of $G$ defined as $G_{x}=\{s\in G:\sigma_{s}(x)=x\}$. Given $s\in G$, we define the set of fixed points of $s$ as $\text{Fix}(s)=\{x\in X:\sigma_{s}(x)=x\}$. It is clear that it is a closed subset of $X$. We define $\text{Per}(\Sigma)=\{x\in X:\sigma_{s}(x)=x\text{ for some }s\in G\setminus\{e\}\}$. If $x$ is in $\text{Per}(\Sigma)$, we say that it is periodic. It is clear that $x$ belongs to $\text{Per}(\Sigma)$ if and only if $G_{x}$ is not trivial. We define $\text{Aper}(\Sigma)$ as $X\setminus\text{Per}(\Sigma)$. If $x$ belongs to $\text{Aper}(\Sigma)$, we say that it is aperiodic or that it has free orbit.
2.2. Representations of $\ell^{1}(\Sigma)$

### v2

# On the involutive Banach algebra associated to topologically free dynamical systems

@@ Abstract.
Given an action $G\curvearrowright X$ of a discrete and countable infinite group $G$ on a compact and Hausdorff space $X$, we regard $\ell^{1}(G\curvearrowright X)$ as the Banach *-algebra crossed product associated to the action. We characterize topological freeness of the action by showing that it is equivalent to every nontrivial closed ideal of $\ell^{1}(G\curvearrowright X)$ intersecting $C(X)$ nontrivially. Most surprisingly, we show that when $G$ is torsion-free and abelian, $\ell^{1}(G\curvearrowright X)$ can detect freeness of $G\curvearrowright X$: indeed, we show that $G\curvearrowright X$ is free if and only if every closed ideal of $\ell^{1}(G\curvearrowright X)$ is self-adjoint, a property that is automatic in $C^{*}$-algebras. We also show with an example that this result does not hold beyond the torsion-free abelian case.
@@ 1. Introduction
The crossed product is a usual construction in the theory of operator algebras that allows one to obtain an algebra from the action of a group on a $C^{*}$-algebra. It is desirable that the crossed product has a Banach algebra structure, and the most common way to do that is by endowing the algebraic skeleton of the algebra with a $C^{*}$-seminorm in order to take the completion of the algebra, which is then a $C^{*}$-algebra. $C^{*}$-crossed products have been studied at length through the years, and are a really important family of examples of $C^{*}$-algebras.
@@
A concrete family of examples comes from an action $G\curvearrowright X$ of a countable infinite discrete group $G$ on a compact Hausdorff space $X$ via homeomorphisms, also called a dynamical system. In this situation, the action of the group on the space gives rise to an action on the algebra $C(X)$ of continuous functions, which is a commutative $C^{*}$-algebra, and from this context one can construct a crossed product endowed with a $C^{*}$-norm. Nevertheless, there are many other types of Banach algebras that can be obtained from the action of a group on a $C^{*}$-algebra, and in the latter years there has been a rise of interest in those. For example, the theory of $L^{p}$-operator algebras has been sparkling in recent years, and an analogue $L^{p}$-operator algebra crossed product can be constructed; see, for example, the survey paper [Gar21] by Gardella. One can also consider the involutive Banach algebra $\ell^{1}(G\curvearrowright X)$, a concrete example whose $C^{*}$-envelope is the usual $C^{*}$-crossed product. In this paper we are interested in the study of this algebra and, in particular, of its ideal structure in relation with the dynamics of the action it arises from. Here, by an ideal we always mean a two-sided ideal. The algebra $\ell^{1}(G\curvearrowright X)$ has a very rich structure. For example, it can have closed non-self-adjoint ideals, which is something that cannot occur in its $C^{*}$-envelope because, in a $C^{*}$-algebra, every closed ideal is self-adjoint. This is one of our main interests, as we are going to see that the existence of closed non-self-adjoint ideals is related to the action not being free. This means that we may detect freeness of the action $G\curvearrowright X$ by looking at the algebra $\ell^{1}(G\curvearrowright X)$, in a way that cannot be done in the $C^{*}$-algebraic crossed product.
For actions of $\mathbb{Z}$, the algebra $\ell^{1}(\mathbb{Z}\curvearrowright X)$ has been studied by de Jeu, Svensson and Tomiyama in [dJST12], by de Jeu and Tomiyama in [dJT12, dJT13, dJT16, dJT18], and by Kishimoto and Tomiyama in [KT18]. More recently, in [JM22], Jauré and Măntoiu studied the $\ell^{1}$-Banach *-subalgebra of a topologically graded $C^{*}$-algebra over a discrete group and, when applying their results to actions of groups, obtained that $\ell^{1}(G\curvearrowright X)$ is inverse closed in its enveloping $C^{*}$-crossed product if the acting group is rigidly symmetric (see [JM22, Proposition 3.1], where the result is stated in the more general framework of partial actions). We are interested in the type of results obtained in [dJST12]. There, it was shown how different dynamical properties are related to analytic-algebraic properties of the ideal structure of $\ell^{1}(\mathbb{Z}\curvearrowright X)$. In particular, [dJST12, Theorem 4.4] states the following: an integer action $\mathbb{Z}\curvearrowright X$ is free (i.e., the action of every element of $G$ apart from the unit has no fixed points) if and only if every closed ideal of $\ell^{1}(\mathbb{Z}\curvearrowright X)$ is self-adjoint.
We are interested in detecting freeness of $G\curvearrowright X$ by looking at the algebra $\ell^{1}(G\curvearrowright X)$ in the broader context of actions of discrete and countable infinite groups. There are several obstructions to extending the arguments used in [dJST12, Theorem 4.4] to actions of more general groups. First of all, the proof uses the existence of a closed non-self-adjoint ideal in $\ell^{1}(\mathbb{Z})$, which is not necessarily true for an arbitrary countable discrete group. It also uses the facts that, for $\mathbb{Z}$-actions, not being free is equivalent to having a finite orbit, and that every nontrivial subgroup of $\mathbb{Z}$ is isomorphic to $\mathbb{Z}$. As it turns out, the statement of [dJST12, Theorem 4.4] does not hold in general for arbitrary groups; see Example 4.1. On the other hand, it can be shown that freeness of the action always implies that every closed ideal is self-adjoint (see Corollary 4.2). In order to prove the other implication we delve into the realm of abelian groups. Imposing some conditions on the group, we obtain the following concrete statement for abelian groups.
@@
Let $G$ be a torsion-free abelian, discrete and countable group. Then, the action $G\curvearrowright X$ is free if and only if every closed ideal of $\ell^{1}(G\curvearrowright X)$ is self-adjoint.
We remark that the analogue of the result above, in the $C^{*}$-algebraic context, does not hold, as closed ideals are always self-adjoint in $C^{*}$-algebras. As such, the $C^{*}$-algebraic crossed product cannot detect freeness of the action the same way that $\ell^{1}(G\curvearrowright X)$ does, and may not even detect freeness at all. This shows a stark contrast between the algebra $\ell^{1}(G\curvearrowright X)$ and its $C^{*}$-envelope.
Different notions of freeness of an action play a key role in this paper. Apart of freeness of the action, the notion of topological freeness will be recurrent (see Definition 3.2). The first landmark that we come across in this paper is the following theorem, in which topological freeness of $G\curvearrowright X$ is characterized through an analytic-algebraic property of $\ell^{1}(G\curvearrowright X)$. In order to make sense of the theorem, we mention that $\ell^{1}(G\curvearrowright X)$ canonically contains the algebra $C(X)$ of continuous functions.
Theorem B.
@@
    For every closed nonzero ideal $I$ of $\ell^{1}(G\curvearrowright X)$, the intersection $I\cap C(X)$ is nonzero.
@@
    For every closed and self-adjoint nonzero ideal of $\ell^{1}(G\curvearrowright X)$, the intersection $I\cap C(X)$ is nonzero.
@@
    The action $G\curvearrowright X$ is topologically free.
@@
We point out the distinction between closed and self-adjoint ideals and just closed ideals done in (1) and (2): as we mentioned before, in this context it is not automatic that every closed ideal is self-adjoint, so the difference is meaningful. This distinction will be recurrent later in the paper. The equivalence between (1) and (3) can actually be seen as an interpretation for $\ell^{1}(G\curvearrowright X)$ of [BK24, Theorem 5.9], which is a result on the more general context of twisted Banach algebra crossed products. However, the proof here is in the spirit of [dJST12, Theorem 4.1], of which it is a generalization for arbitrary discrete and countable (infinite) groups.
@@
Theorem B proves to be an extremely useful tool. Assuming topological freeness of $G\curvearrowright X$, it allows us to easily see how different dynamical properties of the action are characterized as analytic-algebraic properties of $\ell^{1}(G\curvearrowright X)$. The dynamical properties we characterize are minimality, topological transitivity and residual topological freeness (Definitions 3.8, 3.11 and 3.14, respectively); see Theorems 3.10, 3.13 and 3.19. The characterizations of minimality and topological transitivity generalize [dJST12, Theorems 4.2, 4.10], which were proved for $\mathbb{Z}$-actions. In these results, topological freeness of the action is not assumed, because for $\mathbb{Z}$-actions it is automatic given the minimality or topological transitivity. The need to assume topological freeness in the context of arbitrary countable discrete groups does not come as a surprise, given that it was necessary in the context of $C^{*}$-algebraic crossed products; see, for instance, [AS94, Theorem 2].
@@
Some precedents show that algebras of type $L^{1}$ retain more information than their $C^{*}$ counterparts. A great example of this phenomenom is Wendel’s theorem (see [Wen51]), which says that two locally compact Hausdorff groups are isomorphic (as topological groups) if and only if their $L^{1}$-group algebras are isometrically isomorphic. This means that $L^{1}$-group algebras retain all the information of the group, something that the $C^{*}$-group algebras do not satisfy as, for example, the group $C^{*}$-algebras $C^{*}(\mathbb{Z}_{4})$ and $C^{*}(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$ are isomorphic but $\mathbb{Z}_{4}$ is not isomorphic to $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ [^1] In this paper, given $n\in\mathbb{N}$, we refer to $\mathbb{Z}/n\mathbb{Z}$ (the integers modulo $n$) as $\mathbb{Z}_{n}$.. For $C^{*}$-algebraic crossed products, the algebra does not contain that much information of the action by itself. But, given two topologically free actions (see Definition 3.2), an isomorphism between their crossed products that preserves a canonical commutative $C^{*}$-subalgebra is equivalent to the actions being continuous orbit equivalent (see [Li18, Theorem 1.2]). More rigidity is seen on $L^{p}$-operator algebra crossed products, for $p\in[1,\infty)\setminus\{2\}$, as for two topologically free actions, an isometric isomorphism between the crossed product automatically preserves the canonical commutative subalgebra of the crossed products. Then, it holds in this context that the crossed products are isometrically isomorphic if and only if the actions are continuous orbit equivalent (see [CGT24, Theorem 6.7]). There are no rigidity results about $\ell^{1}(G\curvearrowright X)$ as of now, but it is clear that all the information of the action retained by its associated $C^{*}$-crossed product is also retained by $\ell^{1}(G\curvearrowright X)$, as the $C^{*}$-crossed product is the envelope of $\ell^{1}(G\curvearrowright X)$. We hope for some of the results that we present as hinting that the $\ell^{1}$-crossed product may be even more rigid than its counterparts described above.
@@
Let us briefly describe the structure of the paper. In Section 2 we explain the context in which we work, introducing the basic definitions and notations we use. In particular, we introduce $\ell^{1}(G\curvearrowright X)$, our main object of study. In Section 3, we do some technical work in order to prove Theorem B. The characterizations of minimality, topological transitivity and residual topological freeness follow from there. In Section 4, we study how to relate non-freeness of the action with the existence of closed non-self-adjoint ideals in $\ell^{1}(G\curvearrowright X)$. We first prove that if the action is free, then every closed ideal of $\ell^{1}(G\curvearrowright X)$ is self-adjoint. Finally, we show how the existence of a nontrivial point-stabilizer subgroup satisfying some conditions implies the existence of a closed non-self-adjoint ideal in $\ell^{1}(G\curvearrowright X)$ (see Proposition 4.5); this result implies Theorem A.
Acknowledgments
@@
The author is deeply thankful to his late advisor Fernando Abadie, whose guidance was essential to make all this work possible.
@@ 2. Definitions and preliminaries
In this section we introduce the basic objects playing a role in this work, and we fix the notation that we use. In what follows, $G$ is a countable infinite discrete group and $X$ is a compact Hausdorff space. Given a subset $Y$ of $X$, we denote $\overline{Y}$ its closure in $X$.
2.1. The algebra $\boldsymbol{\ell^{1}(G\curvearrowright X)}$
Let $G$ act by homeomorphisms on $X$. For every $s\in G$, we denote by $\sigma_{s}:X\to X$ the homeomorphism associated with the action of $s$ in $X$. We denote this action by $G\curvearrowright X$. We denote by $e$ the identity element of $G$, and we denote by $C(X)=\{f:X\to\mathbb{C};f\text{ is continuous}\}$ the $C^{*}$-algebra of continuous functions on $X$ endowed with the supremum norm $\lVert\>\cdot\>\rVert_{\infty}$. The action of $G$ on $X$ induces an action of $G$ on $C(X)$, that we denote by $\alpha:G\to\text{Aut}(C(X))$, given by
@@ 2.1. The algebra $\boldsymbol{\ell^{1}(G\curvearrowright X)}$
for every $s\in G$, every $f\in C(X)$ and every $x\in X$.
@@
$\ell^{1}(G\curvearrowright X)=\left\{f:G\to C(X):\sum_{s\in G}\lVert f(s)\rVert_{\infty}<\infty\right\}.$
Set $\lVert f\rVert=\sum_{s\in G}\lVert f(s)\rVert_{\infty}$ for every $f\in\ell^{1}(G\curvearrowright X)$. Then $\ell^{1}(G\curvearrowright X)$, with pointwise sum and scalar multiplication, is a Banach space. We endow $\ell^{1}(G\curvearrowright X)$ with the following product: given $f,g\in\ell^{1}(G\curvearrowright X)$, we set
$(fg)(s)=\sum_{t\in G}f(t)\alpha_{t}(g(t^{-1}s))\text{ for every }s\in G.$
We also define an involution $\ell^{1}(G\curvearrowright X)\to\ell^{1}(G\curvearrowright X)$ in the following way: given $f\in\ell^{1}(G\curvearrowright X)$, we set
@@
With these product and involution we have that $\ell^{1}(G\curvearrowright X)$ is a Banach *-algebra. We can canonically embed $C(X)$ in $\ell^{1}(G\curvearrowright X)$ by identifying it (isometrically) with the $\ell^{1}$-functions $G\to C(X)$ supported on the unit $e\in G$.
We now describe a useful way to work with the elements of $\ell^{1}(G\curvearrowright X)$. For every $s\in G$, we define $\delta_{s}\in\ell^{1}(G\curvearrowright X)$ as:
[…]
where $1$ denotes the function taking constant value $1$. With this notation, we can describe an element $f\in\ell^{1}(G\curvearrowright X)$ as:
$f=\sum_{s\in G}f(s)\delta_{s}=\sum_{s\in G}f_{s}\delta_{s},$
where $f_{s}$ denotes $f(s)\in C(X)$ for every $s\in G$. We usually use this description of the elements of $\ell^{1}(G\curvearrowright X)$ without explicit mention. With this notation, we have that
$(f_{s}\delta_{s})(g_{t}\delta_{t})=f_{s}\alpha_{s}(g_{t})\delta_{st}$
for every $s,t\in G$, and every $f_{s},g_{t}\in C(X)$. Also, with this notation we have that the involution is given by the rule
$(f_{s}\delta_{s})^{*}=\alpha_{s}^{-1}(\overline{f_{s}})\delta_{s^{-1}}\text{ for every }s\in G\text{ and }f_{s}\in C(X).$
@@
We define the canonical projection $E:\ell^{1}(G\curvearrowright X)\to C(X)$ as the map given by $E(f)=f(e)$ for every $f\in\ell^{1}(G\curvearrowright X)$.
We denote by $C^{*}(G\curvearrowright X)$ the (universal) $C^{*}$-envelope of $\ell^{1}(G\curvearrowright X)$, which is the universal crossed product $C^{*}$-algebra associated to $G\curvearrowright X$.
We briefly give some definitions related to the action $G\curvearrowright X$. Given $x\in X$, we denote by $G_{x}$ the stabilizer subgroup of $x$ as the subgroup of $G$ defined as $G_{x}=\{s\in G:\sigma_{s}(x)=x\}$. Given $s\in G$, we define the set of fixed points of $s$ as $\text{Fix}(s)=\{x\in X:\sigma_{s}(x)=x\}$. It is clear that it is a closed subset of $X$. We define $\text{Per}(G\curvearrowright X)=\{x\in X:\sigma_{s}(x)=x\text{ for some }s\in G\setminus\{e\}\}$. If $x$ is in $\text{Per}(G\curvearrowright X)$, we say that it is periodic. It is clear that $x$ belongs to $\text{Per}(G\curvearrowright X)$ if and only if $G_{x}$ is not trivial. We define $\text{Aper}(G\curvearrowright X)$ as $X\setminus\text{Per}(G\curvearrowright X)$. If $x$ belongs to $\text{Aper}(G\curvearrowright X)$, we say that it is aperiodic or that it has free orbit.
2.2. Representations of $\boldsymbol{\ell^{1}(G\curvearrowright X)}$